OpenStudy (anonymous):

can you solve this useing quadratic formula?? show me! x^2-11x+8=0

OpenStudy (anonymous):

x = \[-b ± \sqrt{b ^{2} - 4ac}\over2a\] a = 1 b = -11 c = 8 Just plug in numbers.

OpenStudy (anonymous):

that looks like an alien lang. to me :/

OpenStudy (anonymous):

Just plug in the numbers for the letters. \[-(-11) ± \sqrt{(-11)^{2} - 4(1)(8)}\over2(1)\] \[11 ± \sqrt{121 - 32}\over 2\] \[11 ± \sqrt{89}\over2\] Now you can just approximate or if you want an exact answer, just leave it like that.

OpenStudy (anonymous):

trhanks

OpenStudy (anonymous):

You're welcome :)